OpenStudy (jacobbirddog):

Simplify. –10 • 2 ÷ (–4) + 8 • 5 A. 35 B. 65 C. 25 D. 45

OpenStudy (anonymous):

d. 45

OpenStudy (jacobbirddog):

why 45

OpenStudy (anonymous):

P.E.M.D.A.S Order of Operations. \[-10 * 2 = -20\] \[8*5=40\] \[-20 \div (-4) = 5\] \[40 + 5 = 45\]

OpenStudy (jacobbirddog):

thank you just making sure you were not a troll :)

OpenStudy (anonymous):

hahaha no problem